What is the concentration of a solution obtained by diluting 2.50 mL of 5.6 M solution of hydrochloric acid to 50.0 mL with distilled water?
Answer #1

Initial Concentration (M1) = 5.6 M

Initial volume (V1) = 2.50 mL

Final Volume after dilution (V2) =2.50 + 50 = 52.5 mL

So; M1V1 = M2V2

Final conc. after dilution (M2) =M1V1 / V2 = (5.6 M x 2.50 mL) / (52.5 mL) = 0.267 M